# Beyond algorithm hyperparameters: on preprocessing hyperparameters and associated pitfalls in machine learning applications

### v1

@@ 4.2.2 Evaluation of a model generated on a subset of the available data
As in Setting I (see Section 3.2.2), it is also possible in Setting II to use only a subset of the available data for model generation (i.e. $\mathcal{D}_{\text{train}}\subset\mathcal{D}$), and reserve the remaining observations exclusively for evaluation (i.e. $\mathcal{D}_{\text{test}}=\mathcal{D}\setminus\mathcal{D}_{\text{train}}$; see Figure 3, model evaluation d; Hosseini et al.,, 2020). This approach essentially corresponds to nested resampling with holdout as the outer resampling method, except that the holdout is permanent, meaning that the prediction model generated on $\mathcal{D}_{\text{train}}$ (equivalent to $\mathcal{D}^{\prime}_{\text{train}}$ in the previous section) serves as the final prediction model. Similar to Setting I, we thus distinguish the two evaluation procedures by referring to them as temporary outer holdout (described in Section 4.2.1) and permanent outer holdout (described here). We also again note that there might be some confusion in the terminology, as a permanent outer holdout combined with a (temporary) inner holdout can, just like its temporary counterpart, also be referred to as a train-validation-test split.
The statements regarding the temporary vs. permanent holdout in Setting I also apply to Setting II: Compared to the temporary outer holdout, the permanent outer holdout does not exhibit a pessimistic bias as it actually evaluates the final prediction model. However, this comes at the cost of not using all available data for model generation. Accordingly, the same recommendation as in Section 3.2.2 applies: a permanent outer holdout should only be employed if the number of observations in $\mathcal{D}$ is sufficiently large or if it is computationally expensive or practically infeasible to repeat the model generation process. Note that the second point is particularly relevant in Setting II due to the increased effort of model generation (Collins et al., 2024a, ).
@@ 5.1 Real-world prediction problem
Our illustration is based on a real-world prediction problem from the COMPANION study (Hodiamont et al.,, 2022). This study aimed to develop a casemix classification for adult palliative care patients in Germany that considers the complexity of each patient’s palliative care situation to assign them to a class reflecting their resource needs. A casemix classification for palliative care patients has been deemed necessary, as the differentiation of patients based on their diagnosis, which corresponds to the current practice in Germany, has been found to be inappropriate for predicting resource needs in the context of palliative care. Despite yielding many important insights, the COMPANION project was ultimately unable to develop a prediction model with sufficient predictive performance, even after exploring various model generation approaches. However, this makes it a good example to illustrate how optimistically biased evaluation procedures can present prediction models in a more favorable light.
@@
To develop a casemix classification that relates patients’ resource needs to the complexity of their palliative care situation, the COMPANION team formulated a prediction problem where each observation represents a patient’s palliative care phase. The outcome $y^{(i)}$, defined as the average cost per day in palliative care phase $i$, serves as an empirical proxy for resource needs in the corresponding phase. The set of features $\boldsymbol{x}^{(i)}$ intended to reflect the palliative care situation of each phase consists of (i) the type of palliative care phase (categorical), (ii) patient age (integer-valued), (iii) two cognitive features (confusion and agitation; both ordinal), (iv) the Australian Karnofsky Performance Score (AKPS; Abernethy et al.,, 2005) that measures the patients’ functional status (ordinal), and (v) the Integrated Palliative care Outcome Scale (IPOS; Murtagh et al.,, 2019), which is a score that is based on 17 ordinal variables covering physical symptoms, psycho-social burden, family needs, and practical problems. Accordingly, the number of features provided to the learning algorithm is $p=6$. All types of data were collected by the clinical staff of participating palliative care teams.
@@
It is important to note that although the study aimed to identify a casemix classification, the continuous nature of the specified outcome variable (i.e. average cost per day) inherently makes the prediction problem a regression task. To ensure that the obtained prediction model still produces classes that are also interpretable and can be implemented in practice, a decision tree approach was chosen (e.g., using the CART algorithm, discussed in Sections 2-4), despite potential limitations on predictive performance. In the resulting decision tree, each terminal node represents a casemix class (defined by the features that capture the complexity of the palliative care situation) and predicts the average cost per day for that class. Notably, decision trees were also used in the casemix classifications developed for palliative care patients in Australia (Eagar et al.,, 2004) and the UK (Murtagh et al.,, 2023), which served as the basis for many decisions in the development of the German casemix classification.
@@ 5.2.1 Overview
Performing steps (i) to (iii) results in a vector $(\widehat{\mathrm{PE}}_{\text{train}},\widehat{\mathrm{PE}}_{\text{new}})$, which includes the prediction error estimates derived from available and new data, respectively. By comparing these estimates, we can determine whether $\widehat{\mathrm{PE}}_{\text{train}}$ correctly reflects the predictive performance of the model or if it is affected by any form of bias. Ideally, $\widehat{\mathrm{PE}}_{\text{train}}$ should be equal to $\widehat{\mathrm{PE}}_{\text{new}}$, indicating that the model evaluation conducted on $\mathcal{D}_{\text{train}}$ yields an unbiased estimate prediction error estimate (although small differences do not necessarily indicate bias, as $\widehat{\mathrm{PE}}_{\text{new}}$ is also an estimate). To ensure that the difference between the two prediction error estimates is not driven by a specific data split, steps (i) to (iii) are repeated 50 times for each analysis setting (using the same 50 splits for each analysis setting). Since we consider 96 analysis settings and 50 repetitions of splitting the initial COMPANION data set, our illustration generates $96\times 50=4800$ vectors of $(\widehat{\mathrm{PE}}_{\text{train}},\widehat{\mathrm{PE}}_{\text{new}})$. Note that each analysis setting may produce 50 different prediction models, as in each repetition, $\mathcal{D}_{\text{train}}$ contains different observations.
@@
The described setup is implemented in the software environment R (R Core Team,, 2022) using the mlr3 package framework (Lang et al.,, 2019). While the COMPANION data set cannot be made publicly available, the R code and the individual prediction error estimates can be found at https://github.com/NiesslC/overoptimistic_trees.
@@
As stated above, we consider a total of 96 analysis settings. These result from a full factorial variation of four factors: two performance measures, two sample sizes, two learning algorithms, and twelve combinations of model generation and evaluation procedures (yielding the total of $2\times 2\times 2\times 12=96$ analysis settings). The two considered sample sizes are (i) $n_{\text{train}}=724$ (the sample size of $\mathcal{D}_{\text{train}}$ after splitting the original data set) and (ii) $n_{\text{train}}=362$ (half of the observations in $\mathcal{D}_{\text{train}}$ being randomly deleted). Note that $\mathcal{D_{\text{new}}}$ is not affected by this variation and still has $n_{\text{new}}=725$ observations. The two performance measures considered in our illustration are the Root Mean Squared Error (RMSE) and the coefficient of determination ($R^{2}$), which are commonly used performance measures and have also been employed to evaluate other decision-tree-based prediction models for palliative care patients (Eagar et al.,, 2004; Murtagh et al.,, 2023; see Supplementary Section B.3 for more information on both performance measures). Note that in each analysis setting, we use the same performance measure for both the model evaluations performed during model generation (i.e. tuning) and the evaluation of the final prediction model. The two learning algorithms and twelve combinations of model generation and evaluation procedures are described in Sections 5.2.2 and 5.2.3, respectively.
@@ 5.2.2 Learning pipeline and HPs
The learning pipeline $\mathcal{I}$ applied in each training process consists of six preprocessing steps, followed by a learning algorithm (see Figure 4 for an overview). While the full learning pipeline actually consists of more preprocessing steps (referred to in Section 5.1 and detailed in Supplementary Section B.2.1), we will for simplicity not further consider them in the illustration, as they are considered as pre-specified (i.e. have no HPs that are relevant for tuning) and are both parameterless and precede the first parameterized preprocessing step in the learning pipeline (i.e. can safely be applied to the full data set).
@@
After applying all preprocessing steps to the data, it is provided to the learning algorithm, which then yields a prediction model (i.e. a decision tree). We consider two learning algorithms: (i) the CART algorithm (introduced in Section 2.2.1; R package rpart; Therneau and Atkinson,, 2022), and (ii) the Conditional Inference Tree algorithm (CIT; R package partykit; Hothorn and Zeileis,, 2015; Hothorn et al.,, 2006; Zeileis et al.,, 2008). As stated in Sections 2.2.1 and 3.1, the CART algorithm builds a decision tree model by partitioning the feature space $\mathcal{X}$ into terminal nodes using a sequence of binary splits. Since we are considering a regression problem, the splitting rules are determined by minimizing the sum of squared errors, and the prediction value $\hat{f}(\boldsymbol{x})$ for each terminal node is the mean of all outcome values (here: costs) in that node (Breiman et al.,, 1984). The CIT algorithm also employs recursive binary partitioning but instead of minimizing a simple loss function that represents node impurity (here: the sum of squared errors), it uses statistical test procedures to find the optimal splits. This approach has the advantage that, unlike the CART algorithm, the CIT algorithm is not affected by selection bias toward features with many possible splits or missing values (Hothorn et al.,, 2006).
@@
All other HPs of CART and CIT are not tuned and, except for one HP, follow the default values from their corresponding implementation in the mlr3 package (Lang et al.,, 2019), which largely align with the defaults of the underlying packages (i.e. rpart and partykit; Foss and Kotthoff,, 2024). The exception is $\lambda_{maxdepth}$, which we set to 4 to align with the COMPANION project, where this value was chosen to ensure that the resulting decision tree model would be useful in clinical practice.
@@ 5.2.3 Model generation and evaluation procedures
We consider twelve different combinations of model generation and evaluation procedures that could be employed in step (ii) of our illustration (see Section 5.2.1) to obtain a prediction model with associated $\widehat{\mathrm{PE}}_{\text{train}}$. They represent an exemplary yet non-exhaustive selection of procedures that are used in ML applications. The twelve combinations are based on five model generation procedures, where for three of them, we apply two different procedures to evaluate the final prediction model, and for the other two, we use three different evaluation procedures (resulting in the total of $3\times 2+2\times 3=12$ combinations).
@@
Model generation name Pre- specified HPs Tuned HPs Tuning procedure Prediction error estimation Data leakage possible
@@
I I-no tuning $\boldsymbol{\lambda}_{P}$, $\boldsymbol{\lambda}_{A}$ Apparent Yes
@@
II II-manual-P $\boldsymbol{\lambda}_{A}$ $\boldsymbol{\lambda}_{P}$ See Figure  4 None Exhaustive search Sequential Apparent Apparent Yes
@@
II II-automated-A $\boldsymbol{\lambda}_{P}$ $\boldsymbol{\lambda}_{A}$ See Figure  4 60 evaluations Random search Joint 10-fold CV Apparent Yes
@@
II II-combined-PA - $\boldsymbol{\lambda}_{P}$, $\boldsymbol{\lambda}_{A}$ II-manual-P for $\boldsymbol{\lambda}_{P}$ and II-automated-A for $\boldsymbol{\lambda}_{A}$ (for each configuration of $\boldsymbol{\lambda}_{P}$) Apparent Yes
@@
II II-automated-PA - $\boldsymbol{\lambda}_{P}$, $\boldsymbol{\lambda}_{A}$ See Figure  4 210 evaluations Random search Joint 10-fold CV Apparent Yes
@@
In this model generation procedure, the preprocessing HPs ($\boldsymbol{\lambda}_{P}$) are tuned, while the algorithm HPs ($\boldsymbol{\lambda}_{A}$) are set to their default values. It aims to represent inexperienced users who either lack the confidence or the programming skills to tune algorithm HPs but manually experiment with different preprocessing options, without realizing that this is a form of HP tuning. As discussed in Sections 4.1.2 and 4.1.3, manual tuning procedures typically differ from automated tuning procedures, which is reflected by the procedure II-manual-P. First, the HPs are tuned sequentially (i.e. each HP is tuned individually, with previously tuned HPs set to their selected values and subsequently tuned HPs set to their default values). Second, during the tuning of each HP, the apparent error is used to estimate the prediction error of each candidate HP configuration. The order in which the HPs are tuned sequentially is $\lambda_{ipos}$, $\lambda_{age}$, $\lambda_{akps}$, $\lambda_{outlier}$, $\lambda_{ca}$ (which reflects a user who first experiments with variations in the features before removing observations, though any other order is also possible). If more than one HP value yields the same prediction error estimate, the first value that was evaluated is selected. Since the preprocessing HPs are tuned sequentially (i.e. one at a time), and only two ($\lambda_{age}$, $\lambda_{akps}$) or four ($\lambda_{ipos}$, $\lambda_{outlier}$, $\lambda_{ca}$) values per HP are available, only 16 ($=2\times 2+4\times 3$) configurations of $\boldsymbol{\lambda}_{P}$ need to be evaluated during tuning. Therefore, no criterion is specified to terminate tuning before all configurations are evaluated.
@@
This model generation procedure represents a standard procedure in many ML applications, where the algorithm HPs $\boldsymbol{\lambda}_{A}$ are selected through automated tuning, while the preprocessing HPs $\boldsymbol{\lambda}_{P}$ are set to their default values (e.g., because users are not aware that they can be tuned). Even when tuning is fully automated, the procedures used in practice are often simple and based on rules of thumb (Bischl et al.,, 2023), which we aim to reflect in our illustration: we employ a random search algorithm, terminate the tuning after 60 evaluations (which corresponds to 30 times the dimension of the search space, as there are 2 HPs in $\boldsymbol{\lambda}_{A}$), and use 10-fold CV for prediction error estimation. The tuning procedure is performed jointly for all HPs, which is the standard practice for automated tuning.
As with the previous model generation procedures, we report both the apparent error and the 10-fold CV error. Note that, since the 10-fold CV error for the selected HP configuration, $\boldsymbol{\lambda}_{A}^{\mathrm{II}}$, has already been calculated during tuning, we use this value as the 10-fold CV error estimate of the final prediction model to avoid performing additional resampling iterations. Similar to the procedure II-manual-P, data leakage is present in both evaluation procedures and may result in optimistically biased prediction error estimates. Specifically, the optimistic bias in the 10-fold CV error would arise from (resampling-induced) overtuning. Since the procedure II-automated-A is fully automated, we additionally estimate the prediction error using nested CV. Here, we use 10 folds for the outer resampling loop and 2 folds for the inner resampling loop (the small number of inner folds saves computation time and we only need to achieve correct HP selection rather than precise error estimation here; this is also recommended by Bischl et al.,, 2023). As discussed in Section 4.2.1, this evaluation procedure is not affected by data leakage.
@@
The final model generation procedure is similar to the procedure II-automated-A described above, except that the set of jointly tuned HPs now also includes the five preprocessing HPs, $\boldsymbol{\lambda}_{P}$, and the number of evaluations is increased to 210. As in II-automated-A, this corresponds to 30 times the dimension of the search space, as there are now 7 tuned HPs. This procedure represents a conceptually simple way to incorporate preprocessing HPs into the tuning process and is recommended by Bischl et al., (2023). However, as noted in Section 4.1.2, integrating preprocessing HPs into an automated tuning procedure requires advanced programming expertise, which may explain why this procedure is not standard practice yet.

### v2

@@ 4.2.2 Evaluation of a model generated on a subset of the available data
As in Setting I (see Section 3.2.2), it is also possible in Setting II to use only a subset of the available data for model generation (i.e. $\mathcal{D}_{\text{train}}\subset\mathcal{D}$) and reserve the remaining observations exclusively for evaluation (i.e. $\mathcal{D}_{\text{test}}=\mathcal{D}\setminus\mathcal{D}_{\text{train}}$; see Figure 3, model evaluation d; [104]). This approach essentially corresponds to nested resampling with holdout as the outer resampling method, except that the holdout is permanent, meaning that the prediction model generated on $\mathcal{D}_{\text{train}}$ (equivalent to $\mathcal{D}^{\prime}_{\text{train}}$ in the previous section) serves as the final prediction model. Similar to Setting I, we thus distinguish the two evaluation procedures by referring to them as temporary outer holdout (described in Section 4.2.1) and permanent outer holdout (described here). We also again note that there might be some confusion in the terminology, as a permanent outer holdout combined with a (temporary) inner holdout can, just like its temporary counterpart, also be referred to as a train-validation-test split.
The statements regarding the temporary vs. permanent holdout in Setting I also apply to Setting II: Compared to the temporary outer holdout, the permanent outer holdout does not exhibit a pessimistic bias as it actually evaluates the final prediction model. However, this comes at the cost of not using all available data for model generation. Accordingly, the same recommendation as in Section 3.2.2 applies: a permanent outer holdout should only be employed if the number of observations in $\mathcal{D}$ is sufficiently large or if it is computationally expensive or practically infeasible to repeat the model generation process. Note that the second point is particularly relevant in Setting II due to the increased effort of model generation [84].
@@ 5.1 Real-world prediction problem
Our illustration is based on a real-world prediction problem from the COMPANION study [99]. This study aimed to develop a casemix classification for adult palliative care patients in Germany that considers the complexity of each patient’s palliative care situation to assign them to a class reflecting their resource needs. A casemix classification for palliative care patients has been deemed necessary, as the differentiation of patients based on their diagnosis, which corresponds to the current practice in Germany, has been found to be inappropriate for predicting resource needs in the context of palliative care. Despite yielding many important insights, the COMPANION project was ultimately unable to develop a prediction model with sufficient predictive performance, even after exploring various model generation approaches. However, this makes it a good example to illustrate how optimistically biased evaluation procedures can present prediction models in a more favorable light.
@@
To develop a casemix classification that relates patients’ resource needs to the complexity of their palliative care situation, the COMPANION team formulated a prediction problem where each observation represents a patient’s palliative care phase. The outcome $y^{(i)}$, defined as the average cost per day in palliative care phase $i$, serves as an empirical proxy for resource needs in the corresponding phase. The set of features $\boldsymbol{x}^{(i)}$ intended to reflect the palliative care situation of each phase consists of (i) the type of palliative care phase (categorical), (ii) patient age (integer-valued), (iii) two cognitive features (confusion and agitation; both ordinal), (iv) the Australia-modified Karnofsky Performance Status (AKPS; [72]) that measures the patients’ functional status (ordinal), and (v) the Integrated Palliative care Outcome Scale (IPOS; [117]), which is a score that is based on 17 ordinal variables covering physical symptoms, psycho-social burden, family needs, and practical problems. Accordingly, the number of features provided to the learning algorithm is $p=6$. All types of data were collected by the clinical staff of participating palliative care teams.
@@
It is important to note that although the study aimed to identify a casemix classification, the continuous nature of the specified outcome variable (i.e. average cost per day) inherently makes the prediction problem a regression task. To ensure that the obtained prediction model still produces classes that are also interpretable and can be implemented in practice, a decision tree approach was chosen (e.g., using the CART algorithm, discussed in Sections 2-4), despite potential limitations on predictive performance. In the resulting decision tree, each terminal node represents a casemix class (defined by the features that capture the complexity of the palliative care situation) and predicts the average cost per day for that class. Notably, decision trees were also used in the casemix classifications developed for palliative care patients in Australia [92] and the UK [116], which served as the basis for many decisions in the development of the German casemix classification.
@@ 5.2.1 Overview
Performing steps (i) to (iii) results in a vector $(\widehat{\mathrm{PE}}_{\text{train}},\widehat{\mathrm{PE}}_{\text{new}})$, which includes the prediction error estimates derived from available and new data, respectively. By comparing these estimates, we can determine whether $\widehat{\mathrm{PE}}_{\text{train}}$ correctly reflects the predictive performance of the model or if it is affected by any form of bias. Ideally, $\widehat{\mathrm{PE}}_{\text{train}}$ should be equal to $\widehat{\mathrm{PE}}_{\text{new}}$, indicating that the model evaluation conducted on $\mathcal{D}_{\text{train}}$ yields an unbiased estimate prediction error estimate (although small differences do not necessarily indicate bias, as $\widehat{\mathrm{PE}}_{\text{new}}$ is also an estimate). To ensure that the difference between the two prediction error estimates is not driven by a specific data split, steps (i) to (iii) are repeated 50 times for each analysis setting (using the same 50 splits for each analysis setting). Since we consider 96 analysis settings and 50 repetitions of splitting the initial COMPANION data set, our illustration generates $96\times 50=4{,}800$ vectors of $(\widehat{\mathrm{PE}}_{\text{train}},\widehat{\mathrm{PE}}_{\text{new}})$. Note that each analysis setting may produce 50 different prediction models, as in each repetition, $\mathcal{D}_{\text{train}}$ contains different observations.
@@
The described setup is implemented in the software environment R [126] using the mlr3 package framework [113]. While the COMPANION data set cannot be made publicly available, the R code and the individual prediction error estimates can be found at https://github.com/NiesslC/overoptimistic_trees.
@@
As stated above, we consider a total of 96 analysis settings. These result from a full factorial variation of four factors: two performance measures, two sample sizes, two learning algorithms, and twelve combinations of model generation and evaluation procedures (yielding the total of $2\times 2\times 2\times 12=96$ analysis settings). The two considered sample sizes are (i) $n_{\text{train}}=724$ (the sample size of $\mathcal{D}_{\text{train}}$ after splitting the original data set) and (ii) $n_{\text{train}}=362$ (half of the observations in $\mathcal{D}_{\text{train}}$ being randomly deleted). Note that $\mathcal{D_{\text{new}}}$ is not affected by this variation and still has $n_{\text{new}}=725$ observations. The two performance measures considered in our illustration are the Root Mean Squared Error (RMSE) and the coefficient of determination ($R^{2}$), which are commonly used performance measures and have also been employed to evaluate other decision-tree-based prediction models for palliative care patients ([92, 116]; see Supplementary Section B.3 for more information on both performance measures). Note that in each analysis setting, we use the same performance measure for both the model evaluations performed during model generation (i.e. tuning) and the evaluation of the final prediction model. The two learning algorithms and twelve combinations of model generation and evaluation procedures are described in Sections 5.2.2 and 5.2.3, respectively.
@@ 5.2.2 Learning pipeline and HPs
The learning pipeline $\mathcal{I}$ applied in each training process consists of six preprocessing steps, followed by a learning algorithm (see Figure 4 for an overview). While the full learning pipeline actually consists of more preprocessing steps (referred to in Section 5.1 and detailed in Supplementary Section B.2.1), we will, for simplicity, not further consider them in the illustration, as they are considered as pre-specified (i.e. have no HPs that are relevant for tuning) and are both parameterless and precede the first parameterized preprocessing step in the learning pipeline (i.e. can safely be applied to the full data set).
@@
After applying all preprocessing steps to the data, it is provided to the learning algorithm, which then yields a prediction model (i.e. a decision tree). We consider two learning algorithms: (i) the CART algorithm (introduced in Section 2.2.1; R package rpart; [134]), and (ii) the Conditional Inference Tree algorithm (CIT; R package partykit; [106, 105, 142]). As stated in Sections 2.2.1 and 3.1, the CART algorithm builds a decision tree model by partitioning the feature space $\mathcal{X}$ into terminal nodes using a sequence of binary splits. Since we are considering a regression problem, the splitting rules are determined by minimizing the sum of squared errors, and the prediction value $\hat{f}(\boldsymbol{x})$ for each terminal node is the mean of all outcome values (here: costs) in that node [81]. The CIT algorithm also employs recursive binary partitioning, but instead of minimizing a simple loss function that represents node impurity (here: the sum of squared errors), it uses statistical test procedures to find the optimal splits. This approach has the advantage that, unlike the CART algorithm, the CIT algorithm is not affected by selection bias toward features with many possible splits or missing values [105].
@@
All other HPs of CART and CIT are not tuned and, except for one HP, follow the default values from their corresponding implementation in the mlr3 package [113], which largely align with the defaults of the underlying packages (i.e. rpart and partykit; [97]). The exception is $\lambda_{maxdepth}$, which we set to 4 to align with the COMPANION project, where this value was chosen to ensure that the resulting decision tree model would be useful in clinical practice.
@@ 5.2.3 Model generation and evaluation procedures
We consider twelve different combinations of model generation and evaluation procedures that could be employed in step (ii) of our illustration (see Section 5.2.1) to obtain a prediction model with associated $\widehat{\mathrm{PE}}_{\text{train}}$. They represent an exemplary yet non-exhaustive selection of procedures that are used in ML applications. The twelve combinations are based on five model generation procedures, where for three of them, we apply two different procedures to evaluate the final prediction model, and for the other two, we use three different evaluation procedures (resulting in a total of $3\times 2+2\times 3=12$ combinations).
@@
Model
generation
name
Pre-
specified
HPs
Tuned
HPs
Tuning procedure
Prediction
error
estimation
Data
leakage
possible
@@
I I-no tuning
$\boldsymbol{\lambda}_{P}$,
$\boldsymbol{\lambda}_{A}$
- - - - - - Apparent Yes
@@ I I-no tuning
II II-manual-P $\boldsymbol{\lambda}_{A}$ $\boldsymbol{\lambda}_{P}$
See
Figure  4
None
Exhaustive
search
Sequential Apparent Apparent Yes
@@ II II-manual-P $\boldsymbol{\lambda}_{A}$ $\boldsymbol{\lambda}_{P}$
II II-automated-A $\boldsymbol{\lambda}_{P}$ $\boldsymbol{\lambda}_{A}$
See
Figure  4
60
evaluations
Random
search
Joint 10-fold CV Apparent Yes
@@ II II-automated-A $\boldsymbol{\lambda}_{P}$ $\boldsymbol{\lambda}_{A}$
II II-combined-PA -
$\boldsymbol{\lambda}_{P}$,
$\boldsymbol{\lambda}_{A}$
II-manual-P for $\boldsymbol{\lambda}_{P}$ and II-automated-A for $\boldsymbol{\lambda}_{A}$
(for each configuration of $\boldsymbol{\lambda}_{P}$)
Apparent Yes
@@ II II-combined-PA -
II II-automated-PA -
$\boldsymbol{\lambda}_{P}$,
$\boldsymbol{\lambda}_{A}$
See
Figure  4
210
evaluations
Random
search
Joint 10-fold CV Apparent Yes
@@ II II-automated-PA -
In this model generation procedure, the preprocessing HPs ($\boldsymbol{\lambda}_{P}$) are tuned, while the algorithm HPs ($\boldsymbol{\lambda}_{A}$) are set to their default values. It aims to represent inexperienced users who either lack the confidence or the programming skills to tune algorithm HPs but manually experiment with different preprocessing options, without realizing that this is a form of HP tuning. As discussed in Sections 4.1.2 and 4.1.3, manual tuning procedures typically differ from automated tuning procedures, which is reflected by the procedure II-manual-P. First, the HPs are tuned sequentially (i.e. each HP is tuned individually, with previously tuned HPs set to their selected values and subsequently tuned HPs set to their default values). Second, during the tuning of each HP, the apparent error is used to estimate the prediction error of each candidate HP configuration. The order in which the HPs are tuned sequentially is $\lambda_{ipos}$, $\lambda_{age}$, $\lambda_{akps}$, $\lambda_{outlier}$, $\lambda_{ca}$ (which reflects a user who first experiments with variations in the features before removing observations, though any other order is also possible). If more than one HP value yields the same prediction error estimate, the first value that was evaluated is selected. Since the preprocessing HPs are tuned sequentially (i.e. one at a time), and only two ($\lambda_{age}$, $\lambda_{akps}$) or four ($\lambda_{ipos}$, $\lambda_{outlier}$, $\lambda_{ca}$) values per HP are available, only 16 ($2\times 2+4\times 3$) configurations of $\boldsymbol{\lambda}_{P}$ need to be evaluated during tuning. Therefore, no criterion is specified to terminate tuning before all configurations are evaluated.
@@
This model generation procedure represents a standard procedure in many ML applications, where the algorithm HPs $\boldsymbol{\lambda}_{A}$ are selected through automated tuning, while the preprocessing HPs $\boldsymbol{\lambda}_{P}$ are set to their default values (e.g., because users are not aware that they can be tuned). Even when tuning is fully automated, the procedures used in practice are often simple and based on rules of thumb [78], which we aim to reflect in our illustration: we employ a random search algorithm, terminate the tuning after 60 evaluations (which corresponds to 30 times the dimension of the search space, as there are 2 HPs in $\boldsymbol{\lambda}_{A}$), and use 10-fold CV for prediction error estimation. The tuning procedure is performed jointly for all HPs, which is the standard practice for automated tuning.
As with the previous model generation procedures, we report both the apparent error and the 10-fold CV error. Note that, since the 10-fold CV error for the selected HP configuration, $\boldsymbol{\lambda}_{A}^{\mathrm{II}}$, has already been calculated during tuning, we use this value as the 10-fold CV error estimate of the final prediction model to avoid performing additional resampling iterations. Similar to the procedure II-manual-P, data leakage is present in both evaluation procedures and may result in optimistically biased prediction error estimates. Specifically, the optimistic bias in the 10-fold CV error would arise from (resampling-induced) overtuning. Since the procedure II-automated-A is fully automated, we additionally estimate the prediction error using nested CV. Here, we use 10 folds for the outer resampling loop and 2 folds for the inner resampling loop (the small number of inner folds saves computation time, and we only need to achieve correct HP selection rather than precise error estimation here; this is also recommended by [78]). As discussed in Section 4.2.1, this evaluation procedure is not affected by data leakage.
@@
The final model generation procedure is similar to the procedure II-automated-A described above, except that the set of jointly tuned HPs now also includes the five preprocessing HPs, $\boldsymbol{\lambda}_{P}$, and the number of evaluations is increased to 210. As in II-automated-A, this corresponds to 30 times the dimension of the search space, as there are now 7 tuned HPs. This procedure represents a conceptually simple way to incorporate preprocessing HPs into the tuning process and is recommended by [78]. However, as noted in Section 4.1.2, integrating preprocessing HPs into an automated tuning procedure requires advanced programming expertise, which may explain why this procedure is not standard practice yet.